OpenStudy (anonymous):

Which set of line segments could create a right triangle? 15, 30, 35 15, 36, 39 15, 20, 29 5, 15, 30

OpenStudy (anonymous):

Just check it for: \[(one side)^2 = (second)^2 + (third)^2\]

OpenStudy (anonymous):

i was thinking a :D

OpenStudy (anonymous):

If that is true then that will be your answer.. :)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\text{hyotenuse will be the largest side}\\~\\ (hypo)^2=a^2+b^2}\end{align}\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{hypo>a~~~~\\ hypo>b}\end{align}\)